Hi, I am having trouble with the following homework question:
Generate a table and a graph for y=5x-2.
Thank you for your help!

Answer by smik(40)

Great question!


To go about this it would be sensible to make a table first:


How do we make a table? We simply choose values for x, plug them into the equation, and pronto - we've got a value for y. For example, we could find the y value for when x is 2 by plugging 2 in for x. So: y=5(2)-2 >> y=10-2 >> y=8. Therefore, when x is equivalent to 3, y is 8. It is nearly always a good thing to include varied values for x, both negative and positive.


After you've created your table simply plot two points from the table of values you've drawn. If you got y=-2 when x=0, and y=13 when x=3, then plot those two. Then all you have to do is to connect them as we're dealing with a linear equation.


The following is a possible table:
XY
-2-12
-1-7
0-2
13
28



And, a graph like this: